OpenStudy (anonymous):

Find dy/dx by implicit differentiation xy+4x+5x^2=3

OpenStudy (anonymous):

Use product rule for first term 1*y+x y'

OpenStudy (anonymous):

x*dy/dx + y + 4 +10x = 0 y + 4 +10x = -x * dy/dx dy/dx = -(y + 4 +10x) / x

OpenStudy (anonymous):

\[xy'+y+4+10x=3\] solve for y'

OpenStudy (anonymous):

oops that is wrong

OpenStudy (anonymous):

\[xy'+y+4+10x=0\] is more like it

OpenStudy (anonymous):

mancub0 has it

OpenStudy (zarkon):

one can also do it the quick way. If we define \[F(x,y)=x\cdot y+4x+5x^2-3\] then \[\frac{dy}{dx}=-\frac{\frac{\partial F}{\partial x}}{\frac{\partial F}{\partial y}}\] \[=-\frac{y+4+10x}{x}\]